OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#13969 - Optical fiber look - Petite couronne
Incident Report for Web Cloud
Resolved
We have a fiber cut between RAM94 and CAC94.

The ticket was opened at Orange due to a problem of defective garter.

The security of NRAs: CHA92, DEF92, NAN92, RUE92, PUT92, LON92, STC92, MOL92, BGN92, OBS92, CVL92, VIR78, VER78, VEL78, CLA92, ROB92, CAC92, RAM94, VIT94, IVR94, VAU92, CHE78, BNY92 has been compromised.

Update(s):

Date: 2015-07-22 09:10:28 UTC
Problem fixed by replacing a breakout cable.
The loop is once again secure.

Date: 2015-07-22 09:10:02 UTC
The fibre operator confims that ithe issue's not on their side. We will wait for Orange to get back to us.

Date: 2015-07-22 09:09:23 UTC
Just opened a ticket with the operator.

Date: 2015-07-22 09:09:10 UTC
We implemented the measures on Monday and confirmed that there was a cut. We will open a ticket with the operator as Orange has still not intervened.

Date: 2015-07-22 09:08:05 UTC
A technician is on the way to fix the problem, OVH side.

Date: 2015-07-22 09:07:22 UTC
Orange confirmed that they're dealing with the request.
Posted Jul 09, 2015 - 18:49 UTC